A water slide launches a child horizontally above a swimming pool. vertical distance from the top of the slide to the launch point is 3.18 m. The point where the child is launched horizontally is 0.67 m above the water. How far from the launch point does the child land in the water?
Answer #1

vertical distance from the top of the slide to the launch point is 3.18 m.

So at the launch point all the gravitational potential energy is converted into kinetic energy.

PE.=KE.=mgh=(1/2)×m×v2

so v=

V=7.894 m/s

The point where the child is launched horizontally is 0.67 m above the water

So

Time to fall 0.67 m vertical.

H=ut+(1/2)gt2

t=

t=0.3697765459 second

So horizontal distance travelled in this time

distance =speed×time

d=7.894×0.37

d= 2.92 meter

Please Upvote if you get it correct.